# User Centric Semantic Communications

### v1

@@ III-B Task Reflection
After task planning, the transmitter should reflect whether the current task plan can fulfill the user’s request. This step is crucial for the overall system because if the transmitter determines that the current plan cannot satisfy the user’s needs, alternative actions can be taken promptly, preventing any negative impact on the system’s performance. There are two scenarios that may lead to the user’s request not being fulfilled. First, the LLM might select the wrong tool during the Tool Selection step. Second, the toolbox may lack sufficient tools, forcing the LLM to choose an inappropriate tool from the available options. This is a practical concern, given the diverse nature of user’s requests.
The transmitter performs the task reflection by asking the LLM to reflect the current task plan. The prompt that guides the LLM to perform task reflection is summarized in Algorithm 3. Below are two examples.
Example 4 (Positive Reflection Result). User’s request: “Is there a traffic jam in the video”? Current task plan : “Video Sampler $\mid$ Vehicle Density Estimation $\mid$ Analysis”. Task reflection: Yes, this task plan can address the user’s request. The Video Sampler step will help to identify the relevant video frames which is crucial in narrowing down the scope of the analysis to the relevant parts of the video. The Tool Selection step, using the Vehicle Density Estimation tool, will extract the necessary information from the selected video frames. This algorithm is particularly suited to the task as it can compute the vehicle density in a particular video frame, which is a direct indicator of traffic congestion. Finally, the Analysis step will summarize the results from the Vehicle Density Estimation, providing a clear and concise output for the user.
Example 5 (Negative Reflection Result). User’s request: “Did an accident happen in the video”? Current task plan : “Video Sampler $\mid$ Vehicle Motion Detection $\mid$ Analysis”. Task reflection: No, the task plan cannot exactly address the user’s request. The user’s request is to determine if an accident happened in the video. While the framework can select frames, detect vehicle motion, and analyze the results, it does not have the capability to specifically identify an accident.
@@
Algorithm 4 Prompt in Frame Selection
1:Input: {User’s Request}, and {Execution Results of Selected Tool}.
2:The {Execution Results of Selected Tool} only provides partial information and cannot fully fulfill the user’s request. Based on {Execution Results of Selected Tool}, you need to select the most relevant frames that may help to fulfill the user’s request.
3:Do not answer user with this tool’s execution result. You should only output the identification numbers (IDs) of the relevant frames.
4:For example, the user wants to know the number of motorcyclists wearing helmet in the whole video. However, the selected tool can only detect the motorcyclists appears in the video. In this case, you should output the frame IDs of frames with motorcyclists.
@@ IV Frame Selection
When none of the task plans can satisfy the user’s request, according to Algorithm 1, the transmitter should select the most relevant video frames from the surveillance video. This frame selection process can be accomplished by leveraging the LLM and specialized tools.
@@ IV-A Tool Selection
Consider the following two representative user’s requests that none of the task plans can satisfy by applying the toolbox in Table I. The first request is: “Did an accident happen in the video?” We cannot satisfy this request as we lack a tool to detect accident; The second request is “How many motorcyclists wearing helmet in the whole video?” Again, we cannot satisfy this request as we lack a tool to detect motorcyclists while wearing helmets, although we do have a tool to detect motorcyclist. In summary, a reason all task plans fail to meet user’s requests is that the tools in the toolbox are not enough. This is a practical case since the user’s request is diverse while the tools are always limited.
Our strategy is, we ask the LLM to extract one key term from the user’s request, and map it to a tool in the toolbox $T$. In this case, this tool can assist us by selecting relevant video frames. To illustrate this, for the request “Did an accident happen in the video?”, if the key term of this request could be “accident”, then we can apply Vehicle Motion Detection to locate the video frames with non-moving vehicles, since there is a high probability that the vehicles involved in an accident are not moving. As shown in Example 2 and Fig. 3, through the two selected video frames with non-moving cars, the user can determine there is an accident in Fig. 3(b); For the request “How many motorcyclists wearing helmet in the whole video?”, if the key term of this request could be “motorcyclist”, then we can apply Object Detection to locate the video frames with motorcyclists. The user can count the number of motorcyclists wearing helmets by themselves through the selected video frames.
@@
Example 6. Tool Name: Vehicle Motion Detection. Labels: Accident, Collision; Tool Name: Vehicle Detection. Labels : sedan, SUV, van, hatchback, MPV, pickup, bus, truck, and estate.
@@ IV-B Overall Process
The Video Sampler is same as that shown in III-A. For the Frame Selection, we input the execution results from the selected tool to the LLM, and ask the LLM to select the frames that may help to fulfill the user’s request. The prompt that guides the LLM to perform frame selection is summarized in Algorithm 4.
@@ V-A Implementation
We implement the user-centric semantic communication system in the context of traffic surveillance videos, using a computer equipped with an NVIDIA RTX 4070 GPU, and a cell phone. First, the computer acts as the transmitter. It connects remotely to OpenAI’s GPT-4 and stores the surveillance videos locally. Additionally, the 8 specialized AI tools in Table I are installed on the computer and utilize its local resources for detection or estimation tasks. Second, both the cell phone and the computer have WeChat installed (an instant messaging application). The user sends their semantic information requests through WeChat on the cell phone, which are received by the computer. Likewise, the computer sends its feedback—text or video frames—back to the user via WeChat.
@@ V-B Evaluation Methodology and Results
Evaluation Datasets. Our dataset consists of 100 traffic surveillance video clips, each lasting roughly 15 seconds. Among them, 47 video clips are sourced from the datasets CityFlow[19] and CityFlow V2[20], and the other 53 video clips are made by us. In addition, our dataset also consists of 700 possible semantic information requests on the traffic surveillance videos, with 300 labeled “Yes” and 400 labeled “No”. This means, the ground truth is that 300 user’s requests can be fulfilled through our system while 400 cannot be satisfied by any of the generated task plans in our system. We denote Req Y for a request labeled “Yes”; Req N for a request labeled “No”; and Req YN for a general request that is either Req Y or Req N. The dataset has been released to[21]. Last, each video clip is associated with 7 semantic information requests in average.
@@
Performance of Task Reflection in Algorithm 1. We test its performance through the 700 requests in our dataset. Specifically, we input the 700 requests into Algorithm 1, generating corresponding task plans. Then, the task reflection assesses the feasibility of these plans. An accurate evaluation of Req Y occurs when the task reflection believes Req Y can be fulfilled by the task plan; if otherwise, the evaluation is inaccurate. An accurate evaluation of Req N occurs when the task reflection consistently believes Req N cannot be fulfilled by the task plans generated during the $\bar{N}$ iterations in Algorithm 1; if otherwise, the evaluation is inaccurate. The accurate ratio of Req Y is thus defined as the ratio of accurate evaluations of Req Y to the total number of Req Y; the accurate ratio of Req N is defined as the ratio of accurate evaluations of Req N to the total number of Req N; and the accurate ratio of Req YN is defined as the ratio of accurate evaluations of both Req Y and Req N, to the total number of all requests.
@@
The evaluation results are shown in Fig. 4. The results show that, when $\bar{N}=3$, the accurate ratio of the three types of requests exceeds 90%. Specifically, for Req Y, the accurate ratio reaches its maximum value when $\bar{N}\geq 3$. This is because trying more tools can indeed increase the task reflection accurate ratio. However, since once the LLM confirms the selection of a tool, the iterations in Algorithm 1 are terminated. Therefore, increasing of $\bar{N}$ beyond 3 does not further improve accurate ratio; For Req N, the results show that increasing $\bar{N}$ will decrease its accurate ratio. This is due to the fact that, when the task reflection process decides a request cannot be fulfilled, it need to check all $\bar{N}$ tools. If it makes a mistake with any of the $\bar{N}$ tools, the task reflection process may incorrectly believe that the request can be fulfilled. Therefore, a larger $\bar{N}$ can increase the probability of such mistakes. Considering the accurate ratio of Req Y and Req N, as a result, the overall accurate ratio Req NY reaches its maximum at $\bar{N}=3$.
Performance of Overall System. We evaluate the overall user centric semantic communication system through the 100 video clips and the associated 700 user’s requests when $\bar{N}=3$ in Algorithm 1. In addition, we invite 18 postgraduates from 5 Universities to test the performance of the overall system. Specifically, for each user’s request, they receive a feedback from the transmitter (in text form or video frame form). Then, we ask them to evaluate if the transmitter’s feedback can fulfill the user’s request. Evaluations are recorded as either “Yes” or “No”. The user success rate in obtaining semantic information is defined as the ratio of “Yes” evaluations to the total number of requests.
We have conducted 3 independent experiments, in which we test the system’s performance over the dataset repeatedly for 3 times. In addition, for each experiment, we invite 3 different groups of postgraduates to evaluate the system’s performance. The average results are shown in Table II. The results show that our system can fulfill 83.90% of the users’ requests overall, with the success rate for Req Y being higher than that for Req N. This is due to the fact that the performance of Frame Selection is not that good. We leave it as a future work to further improve the performance.
TABLE II: User Success Rate in Obtaining Semantic Information
@@
We also compute the resource savings achieved by our system in terms of frame count (i.e., the number of video frames) and data size, compared with the traditional system in which the whole videos are transmitted to the user. A significant reduction in frame count and data size can substantially lower the transmission bandwidth required for wireless communications. In addition, for the transmission of the video frames, we do not apply any compression techniques for further processing.
From Table III, the frame count is reduced significantly. For the data size, although it is reduced by around 66.33%, the reduction ratio is lower than that of the frame count. A main reason is, although a video clip consists of many video frames where each video frame is with a large data size, advanced compression techniques significantly reduce the video’s overall size. Consequently, the compressed video is much smaller than the combined size of all its video frames. However, when individual video frames are extracted, they retain their large data sizes, resulting in a substantial data size. Thus, the date size reduction ratio is lower than the frame count ratio. We plan to address this problem in a further work.

### v2

@@ III-B Task Reflection
After task planning, the transmitter should assess whether the current task plan can fulfill the receiver’s request. This step is crucial for ensuring the system’s performance. If the transmitter determines that the current plan cannot meet the receiver’s needs, it can take alternative actions to avoid system inefficiencies. There are two scenarios that may lead to the receiver’s request not being fulfilled. First, the LLM might select the wrong tool during the Tool Selection step. Second, the toolbox may lack sufficient tools, forcing the LLM to choose an inappropriate tool from the available options. This is a practical limitation, given the diverse nature of receiver’s requests.
The transmitter performs the task reflection by instructing the LLM to reflect the current task plan. The prompt that guides the LLM to perform task reflection is summarized in Algorithm 3. Below are two examples.
Example 4 (Positive Reflection Result). Receiver’s request: “Is there a traffic jam in the video”? Current task plan : “Video Sampler $\mid$ Vehicle Density Estimation $\mid$ Analysis”. Task reflection: Yes, this task plan can address the receiver’s request. The Video Sampler step will help to identify the relevant video frames which is crucial in narrowing down the scope of the analysis to the relevant parts of the video. The Tool Selection step, using the Vehicle Density Estimation tool, will extract the necessary information from the selected video frames. This algorithm is particularly suited to the task as it can compute the vehicle density in a particular video frame, which is a direct indicator of traffic congestion. Finally, the Analysis step will summarize the results from the Vehicle Density Estimation, providing a clear and concise output for the receiver.
Example 5 (Negative Reflection Result). Receiver’s request: “Did an accident happen in the video”? Current task plan : “Video Sampler $\mid$ Vehicle Motion Detection $\mid$ Analysis”. Task reflection: No, the task plan cannot exactly address the receiver’s request. The receiver’s request is to determine if an accident happened in the video. While the framework can select frames, detect vehicle motion, and analyze the results, it does not have the capability to specifically identify an accident.
@@ IV Frame Selection
When none of the task plans can satisfy the receiver’s request, according to Algorithm 1, the transmitter should select the most relevant video frames from the surveillance video. This frame selection process can be accomplished by leveraging the LLM and specialized tools.
@@ IV-A Tool Selection
Consider the following two representative receiver’s requests that none of the task plans can satisfy by applying the toolbox in Table I. The first request is: “Did an accident happen in the video?” We cannot satisfy this request as we lack a tool to detect accident; The second request is “How many motorcyclists wearing helmet in the whole video?” Again, we cannot satisfy this request as we lack a tool to detect motorcyclists while wearing helmets, although we do have a tool to detect motorcycle. In summary, a reason all task plans fail to meet receiver’s requests is that the tools in the toolbox are not enough. This is a practical case since the receiver’s request is diverse while the tools are always limited.
Our strategy is, we ask the LLM to extract one key term from the receiver’s request, and map it to a tool in the toolbox $T$. In this case, this tool can assist us by selecting relevant video frames. To illustrate this, for the request “Did an accident happen in the video?”, if the key term of this request could be “accident”, then we can apply Vehicle Motion Detection to locate the video frames with non-moving vehicles, since there is a high probability that the vehicles involved in an accident are not moving. As shown in Example 2 and Fig. 3, through the two selected video frames with non-moving cars, the receiver can determine there is an accident in Fig. 3(b); For the request “How many motorcyclists wearing helmet in the whole video?”, if the key term of this request could be “motorcycle”, then we can apply Object Detection to locate the video frames with motorcycles. The receiver can count the number of motorcyclists wearing helmets by themselves through the selected video frames.
@@
Example 6. Tool name: Vehicle Motion Detection. Labels: accident, and collision; Tool name: Vehicle Detection. Labels : sedan, SUV, van, hatchback, MPV, pickup, bus, truck, and estate.
@@ IV-B Overall Process
The Video Sampler is same as that shown in III-A. For the Frame Selection, we input the execution results from the selected tool to the LLM, and ask the LLM to select the frames that may help to fulfill the receiver’s request. The prompt that guides the LLM to perform frame selection is summarized in Algorithm 4.
Algorithm 4 Prompt in Frame Selection
1:Input: {Receiver’s Request}, and {Execution Results of Selected Tool}.
2:The {Execution Results of Selected Tool} only provides partial information and cannot fully fulfill the receiver’s request. Based on {Execution Results of Selected Tool}, you need to select the most relevant frames that may help to fulfill the receiver’s request.
3:Do not answer receiver with this tool’s execution result. You should only output the identification numbers (IDs) of the relevant frames.
4:For example, the receiver wants to know the number of motorcyclists wearing helmet in the whole video. However, the selected tool can only detect the motorcycles appearing in the video. In this case, you should output the frame IDs containing motorcycles.
@@ V-A Implementation
We implement the receiver-centric generative semantic communication system in the context of traffic surveillance videos, using a computer equipped with an NVIDIA RTX 4070 GPU, and a cell phone. First, the computer acts as the transmitter. It connects remotely to OpenAI’s GPT-4 and stores the surveillance videos locally. Additionally, the 8 specialized AI tools in Table I are installed on the computer and utilize its local resources for detection or estimation tasks. Second, the cell phone acts as the receiver. Both the cell phone and the computer have WeChat installed (an instant messaging application). The cell phone sends its semantic information requests through WeChat, which are received by the computer. Likewise, the computer sends its feedback—text or video frames—back to the cell phone via WeChat.
@@ V-B Evaluation Methodology and Results
Evaluation Datasets. Our dataset consists of 100 traffic surveillance video clips, each lasting roughly 15 seconds. Among them, 47 video clips are sourced from the datasets CityFlow[19] and CityFlow V2[20], and the other 53 video clips are made by us. In addition, our dataset also consists of 700 possible semantic information requests on the traffic surveillance videos, with 300 labeled “Yes” and 400 labeled “No”. This means, the ground truth is that 300 receiver’s requests can be fulfilled through our system while 400 cannot be satisfied by any of the generated task plans in our system. For clarity, we denote Req Y for a request labeled “Yes”; Req N for a request labeled “No”; and Req YN for a general request that is either Req Y or Req N. Last, each video clip is associated with 7 semantic information requests in average. The dataset has been made available in[21].
@@
Performance of Task Reflection in Algorithm 1. We test the performance of task reflection through the 700 requests in our dataset. Specifically, we input the 700 requests into Algorithm 1, generating corresponding task plans. Then, the task reflection assesses the feasibility of these plans. An accurate evaluation of Req Y occurs when the task reflection believes Req Y can be fulfilled by the task plan; otherwise, the evaluation is inaccurate. Similarly, an accurate evaluation of Req N occurs when the task reflection consistently believes Req N cannot be fulfilled by the task plans generated during the $\bar{N}$ iterations in Algorithm 1; otherwise, the evaluation is inaccurate. The accurate ratio of Req Y is thus defined as the ratio of accurate evaluations of Req Y to the total number of Req Y; the accurate ratio of Req N is defined as the ratio of accurate evaluations of Req N to the total number of Req N; and the accurate ratio of Req YN is defined as the ratio of accurate evaluations of both Req Y and Req N, to the total number of all requests.
@@
The evaluation results in Fig. 4 show that, when $\bar{N}=3$, the accurate ratios for all request types exceed 90%. Specifically, for Req Y, the accurate ratio reaches its maximum value when $\bar{N}\geq 3$. This is because trying more tools can indeed increase the task reflection accurate ratio. However, since once the LLM confirms the selection of a tool, the iterations in Algorithm 1 are terminated. Therefore, increasing of $\bar{N}$ beyond 3 does not further improve accurate ratio; For Req N, the results show that increasing $\bar{N}$ will decrease its accurate ratio. This is due to the fact that, when the task reflection process decides a request cannot be fulfilled, it need to check all $\bar{N}$ tools. If it makes a mistake with any of the $\bar{N}$ tools, the task reflection process may incorrectly believe that the request can be fulfilled. Therefore, a larger $\bar{N}$ can increase the probability of such mistakes. Balancing the accuracy for both Req Y and Req N, the overall accuracy ratio for Req YN reaches its maximum at $\bar{N}=3$.
TABLE II: Receiver Success Rate in Obtaining Semantic Information
@@
Performance of Overall System. We evaluate the overall receiver-centric generative semantic communication system through the 100 video clips and the associated 700 receiver’s requests when $\bar{N}=3$ in Algorithm 1. In addition, we invite 18 postgraduates from 5 Universities to test the performance of the overall system. Specifically, for each receiver’s request, participants received feedback from the transmitter (either in text or video frame form) and rated whether the feedback successfully addressed the request, recorded as either “Yes” or “No”. The receiver success rate in obtaining semantic information is defined as the percentage of “Yes” responses across all requests.
We have conducted 3 independent experiments, in which we test the system’s performance over the dataset repeatedly for 3 times. In addition, for each experiment, we invite 3 different groups of postgraduates to evaluate the system’s performance. The average results are shown in Table II. The results show that our system can fulfill 83.90% of the receivers’ requests overall. The success rate for Req Y is higher than that for Req N, largely due to limitations in the Frame Selection process. Improving performance of this process will be a focus in future work.
We also compute resource savings achieved by our system in terms of transmitted frame count (i.e., the number of video frames) and data size, compared with the traditional system where the whole videos are transmitted to the receiver for semantic information extraction. A significant reduction in frame count and data size can substantially lower the transmission bandwidth required for wireless communications. Note that we did not apply any additional compression techniques to the video frames during this process.
From Table III, the frame count is reduced significantly. For the data size, although it is reduced by around 66.33%, the reduction ratio is lower than that of the frame count. A main reason is, although a video clip consists of many video frames where each video frame is with a large data size, advanced compression techniques significantly reduce the video’s overall size. Consequently, the compressed video is much smaller than the combined size of all its video frames. However, when individual video frames are extracted, they retain their large data sizes, resulting in a substantial data size. Thus, the data size reduction ratio is lower than the frame count reduction ratio. We plan to address this problem in a future work.